Find the number x that makes the equation correct: 3^x = 9^2 * 3/27^2
Answer by stanbon(75887)

3^x = 9^2 * 3/27^2
-----
3^x = (3^2)^2 * 3/(3^3)^2
----
3^x = 3^4
----
3^x = 3^4
3^x = 3^-1
------
x = -1
--------------
Cheers,
Stan H.
-------------